When a battery is connected to the plates of a 3.00-µF capacitor, it stores a charge of 27.0 µC. What is the voltage of the battery?


1
Expert's answer
2021-05-14T13:57:02-0400


Q=27"\mu C" = "27 \times 10^{-6}" C

C= -3.0 "\mu F" = -3.0 "\times 10^{-6}" F


V= voltage of the battery, C= system capacitance , Q= charge stored by the capacitor


"Q= CV"


"V=\frac{Q}{C}= \frac{27\times 10^{-6} C}{3.0\times 10^{-6}F}"


V=9.0V